Don't understand how to find the prime factorization of 924.
Answer by Earlsdon(6294)

Since 924 is an even number, it will divide by 2, so 2 is a prime factor.
924 = 2*462 Do the same again to the factor 462.
924 = 2*2*231 Now 231 is divisible by 3 , another prime factor, so divide it by 3.
924 = 2*2*3*77 and 77 is divisible by 7, another prime factor so divide by 7.
924 = 2*2*3*7*11 These are all of the prime factors of 924 and you can check by multiplying these prime factors together to get 924.
A prime factor of a number is just a factor of that number that is also a prime number.